# Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination

### v1

@@ Abstract
High-precision spectroscopy of simple atomic systems can be used to advance the theory of atomic energy levels but can also serve as a sensitive probe of nuclear charge radii. For this last purpose, we report an improved measurement of the transition frequency in 4He with 48 Hz uncertainty (0.25 ppt), using a Bose–Einstein condensed sample confined in a magic-wavelength optical dipole trap. A systematic Doppler shift from condensate motion is suppressed by time-resolved ion detection, and the transition frequency is calibrated via a White Rabbit link to a remote active hydrogen maser clock. Combined with previous 3He measurements and improved theory, we obtain the most precise determination to date of the charge-radius difference between the alpha and helion particles of of . This is consistent with other recent determinations and confirms that the current discrepancy between QED theory and experimentally observed ionization energies of excited states in helium is not apparent in the isotope shift.
@@
The spectroscopy laser (bandwidth ) is transfer-locked to an ultrastable laser via an optical frequency comb [30]. The frequency comb and counters used for metrology are referenced via a 110 km long White Rabbit fiber link to a hydrogen maser located at VSL, the national metrology institute of the Netherlands. This link reaches a stability in the low after of averaging, with a noise floor of .
@@ Acknowledgements.
- Tiesinga and Mohr [2023] E. Tiesinga and P. J. Mohr, Tests of fundamental physics, in Springer Handbook of Atomic, Molecular, and Optical Physics, edited by G. W. F. Drake (Springer International Publishing, Cham, 2023) pp. 433–447.
- Morgner et al. [2023] J. Morgner, B. Tu, C. M. König, T. Sailer, F. Heisse, H. Bekker, B. Sikora, C. Lyu, V. A. Yerokhin, Z. Harman, J. R. Crespo López-Urrutia, C. H. Keitel, S. Sturm, and K. Blaum, Stringent test of QED with hydrogen-like tin, Nature 622, 53 (2023).
@@
- Adkins et al. [2022] G. Adkins, D. Cassidy, and J. Pérez-Ríos, Precision spectroscopy of positronium: Testing bound-state qed theory and the search for physics beyond the standard model, Physics Reports 975, 1 (2022), precision spectroscopy of positronium: testing bound-state QED theory and the search for physics beyond the Standard Model.
- Delaunay et al. [2023] C. Delaunay, J.-P. Karr, T. Kitahara, J. C. J. Koelemeij, Y. Soreq, and J. Zupan, Self-consistent extraction of spectroscopic bounds on light new physics, Phys. Rev. Lett. 130, 121801 (2023).
- Tiesinga et al. [2021] E. Tiesinga, P. J. Mohr, D. B. Newell, and B. N. Taylor, Codata recommended values of the fundamental physical constants: 2018, Rev. Mod. Phys. 93, 025010 (2021).
- Beyer et al. [2017] A. Beyer, L. Maisenbacher, A. Matveev, R. Pohl, K. Khabarova, A. Grinin, T. Lamour, D. C. Yost, T. W. Hänsch, N. Kolachevsky, and T. Udem, The Rydberg constant and proton size from atomic hydrogen, Science 358, 79 (2017).
- Fleurbaey et al. [2018] H. Fleurbaey, S. Galtier, S. Thomas, M. Bonnaud, L. Julien, F. Biraben, F. Nez, M. Abgrall, and J. Guéna, New measurement of the transition frequency of hydrogen: Contribution to the proton charge radius puzzle, Phys. Rev. Lett. 120, 183001 (2018).
- Bezginov et al. [2019] N. Bezginov, T. Valdez, M. Horbatsch, A. Marsman, A. C. Vutha, and E. A. Hessels, A measurement of the atomic hydrogen lamb shift and the proton charge radius, Science 365, 1007 (2019).
- Pohl et al. [2010] R. Pohl, A. Antognini, F. Nez, F. D. Amaro, F. Biraben, J. M. R. Cardoso, D. S. Covita, A. Dax, S. Dhawan, L. M. P. Fernandes, A. Giesen, T. Graf, T. W. Hänsch, P. Indelicato, L. Julien, C.-Y. Kao, P. Knowles, E.-O. L. Bigot, Y.-W. Liu, J. A. M. Lopes, L. Ludhova, C. M. B. Monteiro, F. Mulhauser, T. Nebel, P. Rabinowitz, J. M. F. dos Santos, L. A. Schaller, K. Schuhmann, C. Schwob, D. Taqqu, J. F. C. A. Veloso, and F. Kottmann, The size of the proton, Nature 466, 213 (2010).
- van der Werf et al. [2025] Y. van der Werf, K. Steinebach, R. Jannin, H. L. Bethlem, and K. S. E. Eikema, Alpha and helion particle charge radius difference determined from quantum-degenerate helium, Science 388, 850 (2025).
- Henson et al. [2022] B. M. Henson, J. A. Ross, K. F. Thomas, C. N. Kuhn, D. K. Shin, S. S. Hodgman, Y.-H. Zhang, L.-Y. Tang, G. W. F. Drake, A. T. Bondy, A. G. Truscott, and K. G. H. Baldwin, Measurement of a helium tune-out frequency: an independent test of quantum electrodynamics, Science 376, 199 (2022).
- Patkóš et al. [2021] V. Patkóš, V. A. Yerokhin, and K. Pachucki, Complete lamb shift of helium triplet states, Phys. Rev. A 103, 042809 (2021).
- Dorrer et al. [1997] C. Dorrer, F. Nez, B. de Beauvoir, L. Julien, and F. Biraben, Accurate measurement of the two-photon transition frequency in helium: New determination of the lamb shift, Phys. Rev. Lett. 78, 3658 (1997).
- Luo et al. [2016] P.-L. Luo, J.-L. Peng, J. Hu, Y. Feng, L.-B. Wang, and J.-T. Shy, Precision frequency measurements of transitions at 588 nm, Phys. Rev. A 94, 062507 (2016).
- Clausen et al. [2021] G. Clausen, P. Jansen, S. Scheidegger, J. A. Agner, H. Schmutz, and F. Merkt, Ionization energy of the metastable state of from rydberg-series extrapolation, Phys. Rev. Lett. 127, 093001 (2021).
- Clausen et al. [2025] G. Clausen, K. Gamlin, J. A. Agner, H. Schmutz, and F. Merkt, Metrology in a two-electron atom: The ionization energy of metastable triplet helium , Phys. Rev. A 111, 012817 (2025).
- Pachucki et al. [2017] K. Pachucki, V. Patkóš, and V. A. Yerokhin, Testing fundamental interactions on the helium atom, Phys. Rev. A 95, 062510 (2017).
- Wen et al. [2025] J.-L. Wen, J.-D. Tang, Y.-N. Lv, Y. R. Sun, C.-L. Zou, J.-F. Dong, and S.-M. Hu, Postselection shifts the transition frequency of helium in an atomic beam, Science Advances 11, eadu9796 (2025).
- Zheng et al. [2017] X. Zheng, Y. R. Sun, J.-J. Chen, W. Jiang, K. Pachucki, and S.-M. Hu, Measurement of the frequency of the transition of , Phys. Rev. Lett. 119, 263002 (2017).
- Pastor et al. [2004] P. C. Pastor, G. Giusfredi, P. D. Natale, G. Hagel, C. de Mauro, and M. Inguscio, Absolute frequency measurements of the atomic helium transitions around 1083 nm, Phys. Rev. Lett. 92, 023001 (2004).
- Cancio Pastor et al. [2012] P. Cancio Pastor, L. Consolino, G. Giusfredi, P. De Natale, M. Inguscio, V. A. Yerokhin, and K. Pachucki, Frequency metrology of helium around 1083 nm and determination of the nuclear charge radius, Phys. Rev. Lett. 108, 143001 (2012).
- Shiner et al. [1995] D. Shiner, R. Dixson, and V. Vedantham, Three-nucleon charge radius: A precise laser determination using 3He, Phys. Rev. Lett. 74, 3553 (1995).
- van Rooij et al. [2011] R. van Rooij, J. S. Borbely, J. Simonet, M. D. Hoogerland, K. S. E. Eikema, R. A. Rozendaal, and W. Vassen, Frequency metrology in quantum degenerate helium: Direct measurement of the transition, Science 333, 196 (2011).
- Schuhmann et al. [2025] K. Schuhmann, L. M. P. Fernandes, F. Nez, M. A. Ahmed, F. D. Amaro, P. Amaro, F. Biraben, T.-L. Chen, D. S. Covita, A. J. Dax, M. Diepold, B. Franke, S. Galtier, A. L. Gouvea, J. Götzfried, T. Graf, T. W. Hänsch, M. Hildebrandt, P. Indelicato, L. Julien, K. Kirch, A. Knecht, F. Kottmann, J. J. Krauth, Y.-W. Liu, J. Machado, C. M. B. Monteiro, F. Mulhauser, B. Naar, T. Nebel, J. M. F. dos Santos, J. P. Santos, C. I. Szabo, D. Taqqu, J. F. C. A. Veloso, A. Voss, B. Weichelt, A. Antognini, R. Pohl, and T. C. Collaboration, The helion charge radius from laser spectroscopy of muonic helium-3 ions, Science 388, 854 (2025).
- Qi et al. [2025] X.-Q. Qi, P.-P. Zhang, Z.-C. Yan, L.-Y. Tang, A.-X. Chen, T.-Y. Shi, and Z.-X. Zhong, Toward resolving the discrepancy in helium-3 and helium-4 nuclear charge radii, Phys. Rev. Res. 7, L022020 (2025).
- Pachucki et al. [2024] K. Pachucki, V. Patkóš, and V. A. Yerokhin, Second-order hyperfine correction to H, D, and energy levels, Phys. Rev. A 110, 062806 (2024).
@@
- Clausen and Merkt [2025] G. Clausen and F. Merkt, Ionization energy of metastable (2 ) and the alpha- and helion-particle charge-radius difference from precision spectroscopy of the rydberg series, Phys. Rev. Lett. 134, 223001 (2025).
- Rengelink et al. [2018] R. J. Rengelink, Y. van der Werf, R. P. M. J. W. Notermans, R. Jannin, K. S. E. Eikema, M. D. Hoogerland, and W. Vassen, Precision spectroscopy of helium in a magic wavelength optical dipole trap, Nature Physics 14, 1132 (2018).
- Jannin et al. [2022] R. Jannin, Y. van der Werf, K. Steinebach, H. L. Bethlem, and K. S. E. Eikema, Pauli blocking of stimulated emission in a degenerate fermi gas, Nature Communications 13, 6479 (2022), published: 29 October 2022.
- Telle et al. [2002] H. Telle, B. Lipphardt, and J. Stenger, Kerr-lens, mode-locked lasers as transfer oscillators for optical frequency measurements, Applied Physics B: Lasers and Optics 74, 1–6 (2002).
- [31] See the supplementary materials.
- Notermans et al. [2014] R. P. M. J. W. Notermans, R. J. Rengelink, K. A. H. van Leeuwen, and W. Vassen, Magic wavelengths for the transition in helium, Phys. Rev. A 90, 052508 (2014).
- Wu et al. [2018] F.-F. Wu, S.-J. Yang, Y.-H. Zhang, J.-Y. Zhang, H.-X. Qiao, T.-Y. Shi, and L.-Y. Tang, Relativistic full-configuration-interaction calculations of magic wavelengths for the 2 transition of helium isotopes, Phys. Rev. A 98, 040501 (2018).
@@
- Moal et al. [2006] S. Moal, M. Portier, J. Kim, J. Dugué, U. D. Rapol, M. Leduc, and C. Cohen-Tannoudji, Accurate determination of the scattering length of metastable helium atoms using dark resonances between atoms and exotic molecules, Phys. Rev. Lett. 96, 023203 (2006).
- Notermans et al. [2016] R. Notermans, R. Rengelink, and W. Vassen, Comparison of spectral linewidths for quantum degenerate bosons and fermions, Physical Review Letters 117, 213001 (2016).
@@
- Müller et al. [1991] M. W. Müller, A. Merz, M. W. Ruf, H. Hotop, W. Meyer, and M. Movre, Experimental and theoretical studies of the Bi-excited collision systems He*(2 3S) He*(2) at thermal and subthermal kinetic energies, Zeitschrift fur Physik D Atoms, Molecules and Clusters 21, 89–112 (1991).
- Pachucki et al. [2025] K. Pachucki, V. Patkóš, and V. A. Yerokhin, QED nuclear recoil effect in helium isotope shift (2025), arXiv:2512.04623 [physics.atom-ph] .
- Huang et al. [2020] Y.-J. Huang, Y.-C. Guan, J.-L. Peng, J.-T. Shy, and L.-B. Wang, Precision laser spectroscopy of the two-photon transition in , Phys. Rev. A 101, 062507 (2020).
- Sick [2014] I. Sick, Zemach moments of and , Phys. Rev. C 90, 064002 (2014).

### v2

@@ Abstract
High-precision spectroscopy of simple atomic systems can be used to advance the theory of atomic energy levels but can also serve as a sensitive probe of nuclear charge radii. For this last purpose, we report an improved measurement of the transition frequency in 4He with 48 Hz uncertainty (0.25 ppt), using a Bose–Einstein condensed sample confined in a magic-wavelength optical dipole trap. A systematic Doppler shift from condensate motion is suppressed by time-resolved ion detection, and the transition frequency is calibrated via a White Rabbit link to a remote active hydrogen maser clock. Combined with previous 3He measurements and improved theory, we obtain the most precise determination to date of the charge-radius difference between the helion and alpha particle of . This is consistent with other recent determinations and confirms that the current discrepancy between QED theory and experimentally observed ionization energies of excited states in helium is not apparent in the isotope shift.
@@
The spectroscopy laser (bandwidth ) is transfer-locked to an ultrastable laser via an optical frequency comb [32]. The frequency comb and counters used for metrology are referenced via a 110 km long White Rabbit fiber link to an active hydrogen maser located at VSL, the national metrology institute of the Netherlands. This link reaches a stability in the low after of averaging, with a noise floor of .
@@ Acknowledgements.
- [1] Note: See the supplementary materials Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
@@
- [2] (2022) Precision spectroscopy of positronium: testing bound-state qed theory and the search for physics beyond the standard model. Physics Reports 975, pp. 1–61. Note: Precision spectroscopy of positronium: testing bound-state QED theory and the search for physics beyond the Standard Model External Links: ISSN 0370-1573, Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [3] (2017) The Rydberg constant and proton size from atomic hydrogen. Science 358 (6359), pp. 79–85. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [4] (2019) A measurement of the atomic hydrogen lamb shift and the proton charge radius. Science 365 (6457), pp. 1007–1012. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [5] (2012-04) Frequency metrology of helium around 1083 nm and determination of the nuclear charge radius. Phys. Rev. Lett. 108, pp. 143001. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [6] (2025-01) Metrology in a two-electron atom: the ionization energy of metastable triplet helium . Phys. Rev. A 111, pp. 012817. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
@@
- [7] (2021-08) Ionization energy of the metastable state of from rydberg-series extrapolation. Phys. Rev. Lett. 127, pp. 093001. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [8] (2025-06) Ionization energy of metastable (2 ) and the alpha- and helion-particle charge-radius difference from precision spectroscopy of the rydberg series. Phys. Rev. Lett. 134, pp. 223001. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [9] (2023-03) Self-consistent extraction of spectroscopic bounds on light new physics. Phys. Rev. Lett. 130, pp. 121801. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [10] (1997-05) Accurate measurement of the two-photon transition frequency in helium: new determination of the lamb shift. Phys. Rev. Lett. 78, pp. 3658–3661. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [11] (2018-05) New measurement of the transition frequency of hydrogen: contribution to the proton charge radius puzzle. Phys. Rev. Lett. 120, pp. 183001. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [12] (2022) Measurement of a helium tune-out frequency: an independent test of quantum electrodynamics. Science 376 (6589), pp. 199–203. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [13] (2020-06) Precision laser spectroscopy of the two-photon transition in . Phys. Rev. A 101, pp. 062507. External Links: Document, Link Cited by: Figure 5.
- [14] (2022) Pauli blocking of stimulated emission in a degenerate fermi gas. Nature Communications 13 (1), pp. 6479. Note: Published: 29 October 2022 External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [15] (2016-12) Precision frequency measurements of transitions at 588 nm. Phys. Rev. A 94, pp. 062507. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
@@
- [16] (2006-01) Accurate determination of the scattering length of metastable helium atoms using dark resonances between atoms and exotic molecules. Phys. Rev. Lett. 96, pp. 023203. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [17] (2023-10) Stringent test of QED with hydrogen-like tin. Nature 622 (7981), pp. 53–57. External Links: ISSN 1476-4687, Link, Document Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
@@
- [18] (1991) Experimental and theoretical studies of the Bi-excited collision systems He*(2 3S) He*(2) at thermal and subthermal kinetic energies. Zeitschrift fur Physik D Atoms, Molecules and Clusters 21 (2), pp. 89–112. External Links: Link, Document Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [19] (2014-11) Magic wavelengths for the transition in helium. Phys. Rev. A 90, pp. 052508. External Links: Document, Link Cited by: Figure 4, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [20] (2016) Comparison of spectral linewidths for quantum degenerate bosons and fermions. Physical Review Letters 117 (21), pp. 213001. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [21] (2017-06) Testing fundamental interactions on the helium atom. Phys. Rev. A 95, pp. 062510. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [22] (2024-12) Second-order hyperfine correction to H, D, and energy levels. Phys. Rev. A 110, pp. 062806. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [23] (2026) QED nuclear recoil effect in helium isotope shift. Phys. Rev. A 113 (1), pp. 012824. Note: External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [24] (2004-01) Absolute frequency measurements of the atomic helium transitions around 1083 nm. Phys. Rev. Lett. 92, pp. 023001. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [25] (2021-04) Complete lamb shift of helium triplet states. Phys. Rev. A 103, pp. 042809. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [26] (2010) The size of the proton. Nature 466, pp. 213–216. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [27] (2025-04) Toward resolving the discrepancy in helium-3 and helium-4 nuclear charge radii. Phys. Rev. Res. 7, pp. L022020. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [28] (2018) Precision spectroscopy of helium in a magic wavelength optical dipole trap. Nature Physics 14, pp. 1132–1137. External Links: Document Cited by: Table 1, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [29] (2025) The helion charge radius from laser spectroscopy of muonic helium-3 ions. Science 388 (6749), pp. 854–858. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [30] (1995-05) Three-nucleon charge radius: a precise laser determination using 3He. Phys. Rev. Lett. 74, pp. 3553–3556. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [31] (2014-12) Zemach moments of and . Phys. Rev. C 90, pp. 064002. External Links: Document, Link Cited by: Figure 5.
- [32] (2002-01) Kerr-lens, mode-locked lasers as transfer oscillators for optical frequency measurements. Applied Physics B: Lasers and Optics 74 (1), pp. 1–6. External Links: Document Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [33] (2021-06) CODATA recommended values of the fundamental physical constants: 2018. Rev. Mod. Phys. 93, pp. 025010. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [34] (2023) Tests of fundamental physics. In Springer Handbook of Atomic, Molecular, and Optical Physics, G. W. F. Drake (Ed.), pp. 433–447. External Links: ISBN 978-3-030-73893-8, Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [35] (2025) Alpha and helion particle charge radius difference determined from quantum-degenerate helium. Science 388 (6749), pp. 850–853. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [36] (2011) Frequency metrology in quantum degenerate helium: direct measurement of the transition. Science 333 (6039), pp. 196–198. External Links: Document, Link Cited by: Figure 5, Table 1, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [37] (2025) Postselection shifts the transition frequency of helium in an atomic beam. Science Advances 11 (9), pp. eadu9796. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [38] (2018-10) Relativistic full-configuration-interaction calculations of magic wavelengths for the 2 transition of helium isotopes. Phys. Rev. A 98, pp. 040501. External Links: Document, Link Cited by: Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.
- [39] (2017-12) Measurement of the frequency of the transition of . Phys. Rev. Lett. 119, pp. 263002. External Links: Document, Link Cited by: Figure 5, Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination.